# Implicit Subgoal Planning with Variational Autoencoders for Long-Horizon Sparse Reward Robotic Tasks

### v1

@@ Abstract
The challenges inherent to long-horizon tasks in robotics persist due to the typical inefficient exploration and sparse rewards in traditional reinforcement learning approaches. To alleviate these challenges, we introduce a novel algorithm, Variational Autoencoder-based Subgoal Inference (VAESI), to accomplish long-horizon tasks through a divide-and-conquer manner. VAESI consists of three components: a Variational Autoencoder (VAE)-based Subgoal Generator, a Hindsight Sampler, and a Value Selector. The VAE-based Subgoal Generator draws inspiration from the human capacity to infer subgoals and reason about the final goal in the context of these subgoals. It is composed of an explicit encoder model, engineered to generate subgoals, and an implicit decoder model, designed to enhance the quality of the generated subgoals by predicting the final goal. Additionally, the Hindsight Sampler selects valid subgoals from an offline dataset to enhance the feasibility of the generated subgoals. The Value Selector utilizes the value function in reinforcement learning to filter the optimal subgoals from subgoal candidates. To validate our method, we conduct several long-horizon tasks in both simulation and the real world, including one locomotion task and three manipulation tasks. The obtained quantitative and qualitative data indicate that our approach achieves promising performance compared to other baseline methods. These experimental results can be seen in the website https://sites.google.com/view/vaesi/home.
@@
Robotic Manipulation; Task Planning; Reinforcement Learning; Motion Control; Autonomous Agents.
@@ I Introduction
Human daily activities often involve performing long-horizon tasks, which are characterized by hierarchical structures and multiple objectives [1, 2]. These tasks typically comprise numerous simpler yet sequential subtasks, the successful execution of which necessitates long-horizon planning and reasoning capabilities [3, 4, 5]. Humans generally excel at performing long-horizon tasks, largely due to our understanding of the subgoal dependency structure inherent to the task of interest. Although recent advances in the field of intelligent control have demonstrated the effectiveness of robots in performing complex tasks through imitation and reinforcement learning [6, 5, 7], state-of-the-art methods are typically limited to completing either short-term tasks with intensive rewards or long-horizon tasks guided by a series of instructions provided by humans [8]. Empowering robots with the ability to reason and plan long-horizon tasks could increase the areas in which robots can assist or even replace humans, further expanding the frontiers of robotic automation.
@@
Current efforts in robotics to tackle long-horizon tasks involve, e.g., imitation of expert trajectories [9, 10, 6], enhancement of goal-oriented exploration [11, 12, 13, 14, 15] and divide-and-conquer tactics [16, 17, 18]. For instance, imitation learning learns effective strategies to generate subgoals by extracting key points from expert demonstrations. Yet, it usually suffers from time-consuming demonstration collection and sub-optimal expert strategies. While enhancing the exploration efficiency of robots is essential for the discovery of effective and feasible subgoals, the exploration of goals remains limited to neighboring states in proximity to the explored region. Given the conspicuous lack of experience with rarely explored state spaces, a significant gap remains in the efficiency requisite for more generalized reinforcement learning applications [19]. Recently, some research has focused on decomposing complex tasks into simpler ones for solving long-horizon problems. However, most existing methods [16, 17, 18] require an explicit specification of the number of subgoals or decision layers, implying that the entire task’s subgoals must be planned prior to execution, such as in the initial planner shown in Fig. 1 Left. Thus, developing a robot capable of solving an extensive range of combinatorial decision problems continues to pose a long-standing challenge.
Figure 1: Left: Subgoal planning approaches. Right: Explicit and implicit models.
To alleviate these challenges, we propose a novel method, Variational Autoencoder-based Subgoal Inference (VAESI) that solves the long horizon problem by first dividing it into several subtasks in an incremental planner manner (as shown in Fig. 1 Left), and then addressing these subtasks sequentially. Our study mainly focuses on the subgoal generation process (without delving into the underlying subgoal approach process). As shown in Fig. 1 Right, humans can infer subgoals given an initial state and a final goal, and at the same time, they can predict the final goal when they are aware of the initial state and certain subgoals. Inspired by this capability, we formulate two models: one for inferring subgoals based on the current state and the final goal, and the other for reconstructing the final goal using the current state and the generated subgoals. The former model explicitly generates the subgoal directly, while the latter one implicitly approximates the subgoal by maximizing the similarity between the final goal and the reconstructed one. As suggested in [20], we anticipate that adding the implicit model may outperform using only an explicit model when accomplishing subgoal generation for long-horizon tasks.
@@
    We first propose a Variational Autoencoder (VAE)-based Subgoal Generator that employs not only an explicit model to infer subgoals, but also, an implicit model to enhance the quality of the generated subgoals.
@@
    We quantify the quality of inferred subgoals by providing two criteria: feasibility and optimality. We employ a Hindsight Sampler to improve the feasibility of the generated subgoals and a Value Selector to facilitate the selection of optimal subgoals.
@@
    We conduct a series of simulations and real-world experiments to evaluate the performance of our algorithm. The results demonstrate that our method outperforms other approaches under the same environmental settings.
The rest of this paper is organized as follows: Sec. II reviews the state-of-the-art in long-horizon tasks; Sec. III introduces the preliminaries to our algorithm; Sec. IV formulates the problem and objective; Sec. V describes the proposed methodology; Sec. VI presents the experimental results; Sec. VII gives conclusions and future work.
@@ II Related Work
Deep reinforcement learning (DRL) can seek an optimal solution for a given task through well-designed reward functions [21]. Researchers have built on the recent success of DRL by extending this technique to long-horizon tasks with varying goals, known as goal-conditioned reinforcement learning (GCRL). In contrast to standard reinforcement learning, GCRL takes goals into consideration, requiring the agent to accomplish the final goal conditioned on the final goal [22]. Current approaches to solving the long-horizon GCRL problem are mainly imitating expert trajectories, increasing the goal-orientated exploration space, and decomposing complex tasks into small ones.
@@ II-A Imitation Learning
Recent advancements [10, 6] in imitation learning reveal the capacity of humans to identify and learn critical points from expert demonstrations or trajectories to accomplish the final goal. For instance, Paul et al. [9] derive a mapping from current states to subgoals using labels directly acquired from expert trajectories. Jin et al. [23] propose a method enabling the robot to learn an objective function from a few sparsely demonstrated keyframes to solve long-horizon tasks. Joey et al. [24] first utilize representation learning with an instruction prediction loss that operates at a high level of abstraction to accelerate imitation learning results. However, since the trajectories generated by human experts cannot cover the entire region of the state space, the trained policies generalize poorly to unseen tasks. In addition, the strategies trained through imitation learning may not be optimal since the demonstration data probably is not optimal and the trained strategies are unlikely to surpass the demonstration policy. Our approach does not require expert demonstrations and instead employs Reinforcement Learning (RL) to explore the optimal policy through trial and error.
@@ II-B Exploration efficiency
Increasing the exploration efficiency of robots is crucial for solving long-horizon tasks. Numerous studies have endeavored to enhance exploration efficiency by enriching the experience replay buffer to incorporate more unknown states. Works such as Skew-Fit [11], RIG [12], and MEGA [13] prioritize exploring low-density or sparse regions, with the intent of maximizing the likelihood of visiting fewer states, which is equivalent to maximizing the entropy of the desired goal distribution. Warde-Farly et al. [25], on the other hand, aim to learn goal-conditional policies and reward functions by maximizing the mutual information between the achieved and final goals. Conversely, Hartikainen et al. [26] chose subgoals that maximize the distance from the initial state, thereby maximizing the entropy of the desired goal distribution. In general, these methods expand the exploration space but continue to depend on uniformly distributed actions and random action noise. Thus, goal exploration remains limited to neighboring states near the explored region. And there persists a significant gap in the efficiency required for more generalized reinforcement learning applications [19]. Our work tackles the exploration challenge by automatically generating subgoals that can be predicted on the way to the long-horizon final goal.
@@ II-C Subgoal orientated goal-conditioned RL
Drawing inspiration from the human tendency to decompose long-horizon tasks into smaller ones, some researchers are exploring the generation of subgoals by selecting them from historical experiences [27, 16, 28] or by training subgoal models [17, 18]. The experience replay buffer, rich in valuable historical data, facilitates the extraction of valid states as subgoals in a simple and efficient manner, while also ensuring their reachability. For instance, RIS [27] hypothesizes that all states during a particular period can be treated as subgoals for long-duration tasks, leading to random sampling from the experience replay buffer as subgoals. The Hindsight Planner [16] selects subgoals from the experience replay buffer and labels them for supervised learning, resulting in the generation of subgoals. Generally, the subgoal generator exhibits more robustness in predicting subgoals for unknown states. PTP [17] breaks down long-horizon tasks into short tasks by pretraining a Conditional Variation Auto-Encoder (CVAE)-based subgoal generator using offline data, and subsequently fine-tunes the action policy by training each short-term task using online data.
However, these methods necessitate the explicit specification of the number of subgoals or decision layers and rely solely on the initial state to predict future subgoals [17]. Therefore, these methods are inapplicable to tasks conditional on multiple goals, i.e., situations where the same initial state corresponds to multiple goals. Additionally, some distance-based methods [17, 18] strive to maximize the distance between the subgoals and the initial state to obtain better subgoals. They typically assume the co-existence of states and goals within the same space. This assumption does not hold true in numerous environments. For example, in the Stack environment, the state comprises the information of blocks and gripper, whereas the goal pertains solely to the goal position of blocks.
Contrary to previous work that infers subgoals exclusively using an explicit model, we restrict our policy search by employing an implicit model and the prior distribution implicitly represented in the offline dataset. Our work also adopts a divide-and-conquer manner to solve long-horizon tasks by utilizing the hierarchical architecture [29, 30, 31, 32], where the top-level focuses on reasoning and decision-making to generate subgoals, and the low-level interacts with the environment to achieve the subgoals sequentially.
@@ III-A Semi-MDP
We formulate the long-horizon task as a Semi-Markov Decision Process (Semi-MDP) [33]. The Semi-MDP is denoted by $\langle\mathcal{S},\mathcal{A},r,\rho_{0},\rho_{g},\mathcal{T},\gamma,\mathcal{G},\mathcal{O},\psi\rangle$, where $\mathcal{S}$ is state space containing all possible state, $\mathcal{A}$ is action space and $r$ is the reward function. We use $\rho_{0}$ and $\rho_{g}$ to denote the distribution of initial state and final goal, respectively. $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function and $\gamma$ is the discount factor. $\mathcal{G}$ is the goal space from which final goals and subgoals can be sampled. $\mathcal{O}:\langle\mathcal{I},\pi,\mathcal{E}\rangle$ is the option consisting of three components, where $\mathcal{I}\in\mathcal{S}$ is the initial state space, $\mathcal{E}\in\mathcal{S}$ is the terminal state space, $\pi$ is the specific policy for current option, respectively.
@@ III-B Soft Actor-Critic
$J(\pi)=\mathbb{E}_{g\sim\rho_{g},\tau\sim\rho_{\tau}}\left[\sum_{t}\gamma^{t}r(s_{t},a_{t},g)+\alpha\mathcal{H}(\pi(\cdot|s_{t},g))\right]$ (2)
@@ IV Problem Formulation
Within the subgoal space $\mathcal{G}$, the majority of subgoals generated by the option policy are unfeasible from the current state, either due to spatial distance or the strict dependency order of subtasks inherent in long-horizon tasks [37]. An option policy $\psi$ is anticipated to generate a subgoal sequence $\hat{g}_{1},\hat{g}_{2},\cdots,\hat{g}_{n-1}$, such that the transition probability from previous subgoal $\hat{g}_{i-1}$ to current subgoal $\hat{g}_{i}$ exceeds 0. Here, $\hat{g}_{i-1}=\phi(s_{e}^{i-1})=\phi(s_{0}^{i})$ and $\hat{g}_{i}=\phi(s_{e}^{i})$. The transition probability is denoted as $\prod_{s=s_{0}^{i},a\sim\pi_{i}}^{s_{e}^{i}}\left[\pi_{i}(s,\hat{g}_{i})\mathcal{T}(s,a)\right]$. Thus, the feasibility of the whole subgoal sequence can be assured.
@@
$J_{o_{i}}(\psi,\pi_{i})=\mathbb{E}_{\hat{g}_{i}\sim\psi,a_{t}\sim\pi_{i}}\left[\sum_{t}\gamma^{t}r(s_{t}^{i},a_{t},\hat{g}_{i})+\alpha\mathcal{H}(\pi(\cdot|s_{t}^{i},\hat{g}))\right]$ (5)
Figure 2: Examples of feasible and optimal subgoals in the AntMaze environment. The subgoals generated on the blue wall are not feasible for ants. The subgoal sequences shown on the green line tend to achieve higher rewards than those indicated by the red curve.
We utilize AntMaze as a case study to demonstrate the concept of feasible and optimal subgoals. As depicted in Fig. 2, within this “U-shape” maze, an 8-degree-of-freedom quadrupedal ant is initialized at a fixed starting position. The objective of the ant is to reach the final goal, which is the blue sphere in the figure, by manipulating each joint. The option policy may infer subgoals located upon the wall, which are unfeasible for the ants due to spatial inaccessibility, thus the subgoals are not achievable. Fig. 2 also demonstrates two sequences of subgoals, represented in red and green, that guide the ant toward the final goal. It is apparent that the green sequence is more advantageous than the red one, as it enables a faster achievement of the final goal, consequently resulting in a higher reward.
In conclusion, we expect the option policy can infer a sequence of subgoals from the initial state and the final goal, which are feasible and optimal.

### v2

# Explicit-Implicit Subgoal Planning for Long-Horizon Tasks with Sparse Reward

@@ Abstract
The challenges inherent in long-horizon tasks in robotics persist due to the typical inefficient exploration and sparse rewards in traditional reinforcement learning approaches. To address these challenges, we have developed a novel algorithm, termed Explicit-Implicit Subgoal Planning (EISP), designed to tackle long-horizon tasks through a divide-and-conquer approach. We utilize two primary criteria, feasibility and optimality, to ensure the quality of the generated subgoals. EISP consists of three components: a hybrid subgoal generator, a hindsight sampler, and a value selector. The hybrid subgoal generator uses an explicit model to infer subgoals and an implicit model to predict the final goal, inspired by way of human thinking that infers subgoals by using the current state and final goal as well as reason about the final goal conditioned on the current state and given subgoals. Additionally, the hindsight sampler selects valid subgoals from an offline dataset to enhance the feasibility of the generated subgoals. While the value selector utilizes the value function in reinforcement learning to filter the optimal subgoals from subgoal candidates. To validate our method, we conduct four long-horizon tasks in both simulation and the real world. The obtained quantitative and qualitative data indicate that our approach achieves promising performance compared to other baseline methods. These experimental results can be seen on the website https://sites.google.com/view/vaesi.
Note to Practitioners
This paper addresses the persistent challenges in executing long-horizon tasks in robotics, which are hindered by inefficient exploration and sparse rewards in traditional reinforcement learning methods. Existing methods typically struggle with the complexity of long-horizon tasks due to their inability to effectively break down these tasks into manageable subgoals. EISP overcomes this limitation by employing a hybrid subgoal generator, a hindsight sampler, and a value selector. The hybrid subgoal generator utilizes an explicit model to infer subgoals and an implicit model to approximate the subgoals by maximizing the similarity between the final goal and the reconstructed goal. EISP has been validated through rigorous testing on four long-horizon tasks in both simulation and real-world environments. The results demonstrate that EISP offers superior performance compared to existing baseline methods. While preliminary results are promising, future research will delve into integrating with advanced exploration methods and explore its application across a broder range of tasks.
@@
Motion control; Learning control systems; Manipulator motion-planning; Motion-planning.
@@ I Introduction
Human daily activities often involve performing long-horizon tasks, which are characterized by hierarchical structures and multiple distinct types of actions [1, 2]. These tasks typically comprise numerous simpler yet sequential subtasks, the successful execution of which necessitates long-horizon planning and reasoning capabilities [3, 4]. Humans generally excel at performing long-horizon tasks, largely due to our understanding of the subgoal dependency structure inherent to the task of interest. Although recent advances in the field of intelligent control have demonstrated the effectiveness of robots in performing complex tasks through imitation and reinforcement learning [5, 4, 6], state-of-the-art methods are typically limited to completing either short-term tasks with intensive rewards or long-horizon tasks guided by a series of instructions provided by humans [7]. Empowering robots with the ability to reason and plan long-horizon tasks could increase the areas in which robots can assist or even replace humans, further expanding the frontiers of robotic automation.
Figure 1: (a) EISP infers multiple subgoals at different task stages. Green and red lines represent possible subgoal sequences. (b) We leverage the explicit model to infer subgoals and the implicit model to predict the final goal.
@@
Current efforts in robotics to tackle long-horizon tasks involve, e.g., imitation of expert trajectories [8, 9, 5], enhancement of goal-oriented exploration [10, 11, 12, 13, 14] and divide-and-conquer tactics [15, 2, 16]. For instance, imitation learning learns effective strategies to generate subgoals by extracting key points from expert demonstrations. Yet, it usually suffers from time-consuming demonstration collection and sub-optimal expert strategies. While enhancing the exploration efficiency of robots is essential for the discovery of effective and feasible subgoals, the exploration of goals remains limited to neighboring states of the explored region. Given the conspicuous lack of experience with rarely explored state spaces, a significant gap remains in the efficiency requisite for more generalized reinforcement learning applications [17]. Recently, some research has focused on decomposing complex tasks into simpler ones for solving long-horizon problems. However, most existing methods [15, 2, 16] cannot simultaneously guarantee the feasibility and optimality of the generated subgoal sequence and often operate under the strong assumption that the state space and the goal space are the same. Thus, developing a robot capable of solving an extensive range of combinatorial decision problems continues to pose a long-standing challenge.
To alleviate these challenges, we propose Explicit-Implicit Subgoal Planning (EISP) that solves the long horizon problem by inferring easy-to-achieve subgoals and then accomplishing each subgoal sequentially. Fig.1 (a) illustrates the potential subgoals that can be generated at different stages from the initial state to the final goal. The subgoal sequences connected by the green line and the red line represent possible approaches to accomplishing the long-horizon task. Because the subgoal sequence linked by the red line may require more time to complete the task, we consider the subgoal sequence connected by the green line is more optimal. Our study mainly focuses on the subgoal generation process (without delving into the underlying subgoal approach process). Inspired by the human approach to solving long-horizon tasks by breaking them down into smaller and manageable components, it is intuitive to employ an explicit model that takes as input the current state and final goal and outputs the subgoals, as demonstrated in previous studies [2, 15]. However, a critical aspect often overlooked is humans’ ability to predict long-term outcomes by identifying short-term subgoals based on the current state. As depicted in Fig.1 (b), we formulate two models: one for inferring subgoals based on the current state and the final goal, and the other for reconstructing the final goal using the current state and the given subgoals. The former model explicitly generates the subgoal, whereas the latter provides a guarantee on the worst-case for the log-likelihood of the subgoal distribution. As suggested in [18], we anticipate that adding the implicit model may outperform using only an explicit model when accomplishing subgoal generation for long-horizon tasks.
@@
    We introduce a Hybrid Subgoal Generator, which employs both explicit and implicit models to generate accurate subgoals by maximizing the evidence lower bound of the final goal.
@@
    We quantify the quality of inferred subgoals by establishing two criteria: feasibility and optimality. We employ a Hindsight Sampler to facilitate the feasibility of the generated subgoals and a Value Selector to facilitate the selection of optimal subgoals.
@@
    We conduct a series of simulations and real-world experiments to evaluate the performance of our algorithm. The results demonstrate that our method outperforms other approaches under comparable environmental conditions.
The rest of this paper is organized as follows: Sec. II reviews the state-of-the-art; Sec. III introduces the preliminaries; Sec. IV formulates the problem; Sec. V describes the proposed methodology; Sec. VI presents the experimental results; Sec. VII gives conclusions and future work.
@@ II Related Work
Deep reinforcement learning (DRL) can seek an optimal solution for a given task through well-designed reward functions. Researchers have built on the recent success of DRL by extending this technique to long-horizon tasks with varying goals, known as goal-conditioned reinforcement learning (GCRL). Current approaches to solving the long-horizon GCRL problem are mainly imitating expert trajectories, increasing the goal-orientated exploration space, and decomposing complex tasks into small ones.
@@ II-A Imitation Learning
Recent advancements [9, 5] in imitation learning reveal the capacity of humans to identify and learn critical points from expert demonstrations or trajectories to accomplish the final goal. For instance, Paul et al. [8] derive a mapping from current states to subgoals using labels directly acquired from expert trajectories. Jin et al. [19] propose a method enabling the robot to learn an objective function from a few sparsely demonstrated keyframes to solve long-horizon tasks. [20] learns dual-arm manipulation tasks from demonstrations generated by a teleoperation system. Joey et al. [21] first utilize representation learning with an instruction prediction loss that operates at a high level of abstraction to accelerate imitation learning results. However, since the trajectories generated by human experts cannot cover the entire region of the state space, the trained policies generalize poorly to unseen tasks. In addition, the strategies trained through imitation learning may not be optimal since the demonstration data probably is not optimal and the trained strategies are unlikely to surpass the demonstration policy. Our approach does not require expert demonstrations and instead employs Reinforcement Learning (RL) to explore the optimal policy through trial and error.
@@ II-B Exploration efficiency
Increasing the exploration efficiency of robots is crucial for solving long-horizon tasks. Numerous studies have endeavored to enhance exploration efficiency by enriching the experience replay buffer to incorporate more unknown states. Works such as Skew-Fit [10], RIG [11], and MEGA [12] prioritize exploring low-density or sparse regions, with the intent of maximizing the likelihood of visiting fewer states, which is equivalent to maximizing the entropy of the desired goal distribution. Warde-Farly et al. [22], on the other hand, aim to learn goal-conditional policies and reward functions by maximizing the mutual information between the achieved and final goals. Conversely, Hartikainen et al. [23] chose subgoals that maximize the distance from the initial state, thereby maximizing the entropy of the desired goal distribution. In general, these methods expand the exploration space but continue to depend on uniformly distributed actions and random action noise. Thus, goal exploration remains limited to neighboring states near the explored region. Our work tackles the exploration challenge by automatically generating subgoals that can be predicted on the way to the long-horizon final goal.
@@ II-C Subgoal orientated goal-conditioned RL
Drawing inspiration from the human tendency to decompose long-horizon tasks into smaller ones, some researchers are exploring the generation of subgoals by selecting them from historical experiences [24, 15, 25] or by training subgoal models [2, 16]. The experience replay buffer, rich in valuable historical data, facilitates the extraction of valid states as subgoals in a simple and efficient manner, while also ensuring their reachability. However, some methods necessitate the explicit specification of the number of subgoals or decision layers and rely solely on the initial state to predict future subgoals [2]. And distance-based methods, like [2, 16] strive to maximize the distance between the subgoals and the initial state to obtain better subgoals. They typically assume the co-existence of states and goals within the same space. This assumption does not hold true in numerous environments. For example, in the Stack environment, the state comprises the information of blocks and gripper, whereas the goal pertains solely to the goal position of blocks.
Recently, new image editing diffusion models like SuSIE [26] and SkillDiffuser [27] have been introduced for subgoal generation. Despite their powerful capabilities, these models depend on pretrained large diffusion models and robot demonstration video data, and are confined to image-based environments. Moreover, the substantial computational resources required for these large models pose a challenge for deployment in real-world robotic tasks.
Contrary to previous work that infers subgoals exclusively using an explicit model, we restrict our policy search by employing an implicit model and the prior distribution implicitly represented in the offline dataset. Our work also adopts a divide-and-conquer manner to solve long-horizon tasks by utilizing the hierarchical architecture [28, 29, 30], where the top-level focuses on reasoning and decision-making to generate subgoals, and the low-level interacts with the environment to achieve the subgoals sequentially.
@@ III-A Semi-MDP
We formulate the long-horizon task as a Semi-Markov Decision Process (Semi-MDP) [31]. More precisely, we extend the general MDP by adding option policy $\psi$ and options $\mathcal{O}$. The Semi-MDP is denoted by $\langle\mathcal{S},\mathcal{A},r,\rho_{0},\rho_{g},\mathcal{T},\gamma,\mathcal{G},\mathcal{O},\psi\rangle$, where $\mathcal{S}$ is state space containing all possible state, $\mathcal{A}$ is action space and $r$ is the reward function. We use $\rho_{0}$ and $\rho_{g}$ to denote the distribution of initial state and final goal, respectively. $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function and $\gamma$ is the discount factor. $\mathcal{G}$ is the goal space from which final goals and subgoals can be sampled. $\mathcal{O}:\langle\mathcal{I},\pi,\mathcal{E}\rangle$ is the option consisting of three components, where $\mathcal{I}\in\mathcal{S}$ is the initial state space, $\mathcal{E}\in\mathcal{S}$ is the terminal state space, $\pi$ is the specific policy for current option, respectively.
@@ III-B Soft Actor-Critic
{dmath}
J(π) = E_g∼ρ_g, τρ_τ [ ∑_t γ^t r(s_t, a_t, g) + αH(π(⋅—s_t,g)) ]
@@ IV Problem Formulation
Within the subgoal space $\mathcal{G}$, the majority of subgoals generated by the option policy are unfeasible, either due to spatial distance or the strict dependency order of subtasks inherent in long-horizon tasks [35]. An option policy $\psi$ is anticipated to generate a subgoal sequence $\hat{g}_{1},\hat{g}_{2},\cdots,\hat{g}_{n-1}$, such that the transition probability from previous subgoal $\hat{g}_{i-1}$ to current subgoal $\hat{g}_{i}$ exceeds 0. Here, the transition probability is denoted as $\rho(s_{0}^{i})\prod_{s=s_{0}^{i},a\sim\pi_{i}}^{s_{e}^{i}}\left[\pi_{i}(s,\hat{g}_{i})\mathcal{T}(s,a)\right]$ and $\rho(s_{0}^{i})$ is the probability of state $s_{0}^{i}$. More details can be found in Appendix -A.
@@
$J_{o_{i}}(\psi,\pi_{i})=\mathbb{E}_{\hat{g}_{i}\sim\psi,a_{t}\sim\pi_{i}}\Big{[}\sum_{t}\gamma^{t}r(s_{t}^{i},a_{t},\hat{g}_{i})+\alpha\mathcal{H}(\pi(\cdot|s_{t}^{i},\hat{g}))\Big{]}$ (3)
Figure 2: Examples of feasible and optimal subgoals in the OpenDrawer environment are depicted.
We employ the OpenDrawer environment as a case study to elucidate the concept of feasible and optimal subgoals. As illustrated in Fig. 2, the ultimate objective is to open the drawer by manipulating the handle. However, a red obstacle block is positioned in front of the drawer, necessitating that the robotic arm first push the block away. This task encompasses two distinct skills: pushing the block and opening the drawer. The option policy infers subgoals within the subgoal space, which can sometimes interfere with the drawer, rendering them unfeasible for the robot. Considering the spatial relationship between the block and the drawer, the subgoal encircled by the yellow line is deemed optimal as it enables the robot to move the block away more efficiently.